Fermi National Laboratory
 

Operation Summary for October 7, 2009 - October 9, 2009

 

Wednesday Oct. 7

The day shift began with Operations monitoring the Tevatron (TeV) store (63.36E30) and Recycler stash (313.3E10), with the Antiproton Source stacking (13.8E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam.

During the day shift, Operations delivered beam to the Pbar Source, NuMI, and MiniBooNE when available. Store 7246 collided until the scheduled termination time of 12:20 PM. Operations put the TeV into HEP shot setup at 12:31 PM and established store 7247 at 2:06 PM with an average initial luminosity (AIL) of 245.89E30.

During the evening shift, Operations delivered beam to the Pbar Source, NuMI, and MiniBooNE when available. Linac output fell from 31mA to 23mA, by tuning the Buncher phase the expert was able to recover to nominal intensities.

Thursday Oct. 8

The midnight shift began with Operations monitoring store 7247 (88.54E30) and the stash (240.4E10), with Pbar stacking (10.1E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations delivered beam to the Pbar Source, NuMI, and MiniBooNE when available. Store 7247 collided until the scheduled termination time of 7:49 AM.

During the day shift, Operations put the TeV into HEP shot setup at 8:00 AM and established store 7248 at 10:12 AM with an average initial luminosity (AIL) of 228.45E30.

During the evening shift, Operations delivered beam to the Pbar Source, NuMI, and MiniBooNE when available.

Friday Oct. 9

The midnight shift began with Operations monitoring store 7248 (71.76E30) and the stash (330.8E10), with Pbar stacking (26.8E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations delivered beam to the Pbar Source, NuMI, and MiniBooNE when available. Store 7248 collided until the scheduled termination time of 2:19 AM. Operations put the TeV into HEP shot setup at 2:31 AM and established store 7249 at 4:21 AM with an average initial luminosity (AIL) of 236.32E30.

The Plans for Friday and the Weekend

The plan for today is a 4 hour Tevatron damper study, along with a CDF access. The plans for the weekend are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/09/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory